allison is saving money to buy a video game system. in the first week, her savings were $8 less then 2/5 of the price of the system. in the second week, she saved 50 cents more than 1/2 the price of the system. she was still $37 short. whats the price of the system?
Answer by mananth(16949)

allison is saving money to buy a video game system. in the first week, her savings were $8 less then 2/5 of the price of the system. in the second week, she saved 50 cents more than 1/2 the price of the system. she was still $37 short. whats the price of the system?
Let system cost $x
First week -- (2/5)x -8
II week-------(1/2)x +0.5
She was$37 short
Add up
savings + short = price of system
(2/5)x-8 + (1/2)x+0.5 +37 =x
(2/5)x +(1/2)x -x= -37-0.5+8
(4x+5x-10x)/10 = -29.5
-x = 10*-29.5
-x=-295
x= $ 295 the price of the system